The line 𝑦=π‘₯βˆ’5 is tangent to the curve 𝑦=π‘₯^2βˆ’9π‘₯βˆ’π‘˜

Find the value of π‘˜.


1
Expert's answer
2022-02-08T10:57:12-0500

The slope of the tangnet line "y=x-5" is "m=1".

"y=x^2-9x-k" , "\frac{dy}{dx} =2x-9=1"


We have "2x-9=1" ,

"2x=10" ,

"x=5".


Then "y=x-5=0".


So, The line "y=x-5" is tangent to the curve "y=x^2-9x-k" at the point "(5, \ 0)" .


"0=5^2-9\cdot 5-k"

"k=-20"



Answer: "k=-20"